# Unlocking State-Tracking in Linear RNNs Through Negative Eigenvalues

### v1

@@ 1 Introduction
Despite these successes, both Transformers and current LRNNs face a fundamental limitation: they struggle to learn how to track the state of even simple finite-state machines from sequences of state-transitions (Deletang et al., 2023). This limitation may impair performance on tasks such as entity tracking in narratives, handling nested structures in code, and other reasoning tasks that can benefit from maintaining and updating an internal state over time (Merrill et al., 2024). Even the simplest state-tracking task, computing the parity of a sequence of bits, cannot be solved by current LRNNs, while non-linear RNNs like LSTM (Hochreiter & Schmidhuber, 1997) and sLSTM (Beck et al., 2024) can solve parity (Merrill et al., 2024). However, in contrast to modern linear RNNs, non-linear RNNs lack an efficient method for parallelizing the training across the sequence length.
Recently, Sarrof et al. (2024) demonstrated that the inability of diagonal LRNNs to solve the parity problem stems from the fact that the eigenvalues of their state-transition matrices are constrained to be positive. Specifically, they proved that finite precision diagonal LRNNs with exclusively positive real eigenvalues, cannot solve the parity problem for sequences of arbitrary length. However, their work did not provide empirical evidence showing that diagonal LRNNs with negative eigenvalues can be successfully trained to overcome this limitation. We prove that the same limitation also affects LRNNs with non-diagonal state-transition matrices, and further prove that complex eigenvalues are necessary to solve the more challenging task of modular counting (when the modulus is not a power of two). Our findings also apply to the GH matrices employed by DeltaNet, as they share the same eigenvalue limitations. To overcome this, we propose a simple yet powerful solution: extend the range of possible eigenvalues from $[0,1]$ to $[-1,1]$. This change enables state-tracking and significantly improves the expressivity of LRNNs without compromising their efficiency and training stability. As illustrated in Figure 1, it allows diagonal LRNNs to learn parity successfully. The code for part of our experiments is available at this link.
@@
    By extending the eigenvalue range, we significantly improve the state-tracking capabilities of LRNNs. We prove that LRNNs with state-transition matrices formed by products of generalized Householder (GH) matrices, each with eigenvalues in the range $[-1,1]$, can learn any regular language (Theorem 4), in some cases with just one layer (Theorem 3). Notably, this range extension allows LRNNs, using just one GH matrix (like DeltaNet), to learn substantially harder tasks, such as the composition of permutations of two elements, compared to diagonal LRNNs.
@@ 2 Related Work
Linear RNNs.  Linear RNNs encompass state-space models and causal, linear attention mechanisms. State-space models, originally used for continuous dynamical systems, inspired LRNN variants like S4 (Gu et al., 2022) and H4 (Fu et al., 2021) (see Tiezzi et al. (2024) for a survey). Recent advancements, such as Mamba (Gu & Dao, 2023; Dao & Gu, 2024), introduced input-dependent gating of the hidden state, significantly improving language modeling performance. Concurrently, linear attention emerged as an alternative to classical softmax attention, with Katharopoulos et al. (2020) demonstrating that causal, linear attention Transformers can be reformulated as RNNs with linear scaling in sequence length. Building on this, Yang et al. (2024a) proposed Gated Linear Attention (GLA), adding a gating mechanism similar to Mamba, while DeltaNet (Yang et al., 2024b) and TTT-Linear (Sun et al., 2024) explored more expressive gating with non-diagonal state-transition matrices. Recent work has combined non-linear and linear RNNs, as seen in xLSTM (Beck et al., 2024), a successor to the traditional LSTM (Hochreiter & Schmidhuber, 1997).
@@
Expressivity Results.  Several studies have explored the expressive power of Transformers and RNNs (see e.g. (Merrill et al., 2020; Strobl et al., 2024; Bhattamishra et al., 2024)). Here, we focus on the ones most relevant to our work. While Hahn (2020) proved that Transformers cannot model periodic languages such as parity and some context-free languages at arbitrary sequence lengths, Liu et al. (2023) demonstrated that Transformers can learn shortcut solutions for solvable finite state automata, though these solutions lack generalizability to arbitrary sequence lengths and perform poorly out-of-distribution. Unlike RNNs, the high parallelizability of Transformers prevents them from learning unsolvable finite state automata (Merrill & Sabharwal, 2023). These findings typically use techniques from algebraic formal language theory (we refer to Liu et al. (2023) for a short tutorial) and circuit complexity, using the log-precision assumption and a number of layers scaling linearly or logarithmically with sequence length. While earlier research established Transformers’ Turing completeness, it relied on either arbitrary precision (Pérez et al., 2021) or arbitrary depth and weight sharing (Giannou et al., 2023). Diagonal LRNNs can simulate any RNN with infinite depth (Gu & Dao, 2023) and approximate regular enough functions when the state dimension grows linearly with sequence length (Orvieto et al., 2024). However, things change when depth and state size are fixed. Merrill et al. (2024) proved that finite-depth diagonal LRNNs, like Transformers, cannot learn unsolvable finite state automata when restricted to log-precision arithmetic. The work by Fan et al. (2024) highlights a similar limitation, while in a finite precision setting, Sarrof et al. (2024) showed that diagonal LRNNs with positive values in the state-transition matrix, while capable of learning all star-free languages, cannot solve even the simple parity problem, a non-star-free language recognizable by a solvable automaton with two states. However, their analysis was limited to the diagonal case and they did not test the benefit of negative eigenvalues in practice. Unlike these works, we also study non-diagonal LRNNs that can still be trained efficiently at large scale.
@@ 3.1 Linear Recurrent Neural Networks (LRNNs)
Table 1: Instances of LRNNs layers in (1), where $\bm{\alpha}_{t}{=}\mathrm{sigmoid}({\bm{W}}_{\alpha}{\bm{x}}_{t})$, $\bm{\Delta}_{t}{=}\mathrm{softplus}({\bm{W}}_{\Delta}{\bm{x}}_{t})$, $\beta_{t}{=}\mathrm{sigmoid}({\bm{w}}_{\beta}{\bm{x}}_{t})$, while ${\bm{q}}_{t},{\bm{k}}_{t}\in\mathbb{R}^{n},{\bm{v}}_{t}\in\mathbb{R}^{d}$ are output of learnable possibly non-linear functions of ${\bm{x}}_{t}$. Also $\psi:\mathbb{R}^{d}\to\mathbb{R}^{d}$ is another learnable function usually containing an MLP and a normalization, while ${\bm{W}}_{1}\in\mathbb{R}^{n\times d}$, ${\bm{W}}_{\Delta}\in\mathbb{R}^{d\times l}$, ${\bm{W}}_{\alpha}\in\mathbb{R}^{n\times l}$, ${\bm{w}}_{\beta}\in\mathbb{R}^{l}$ and ${\bm{w}}_{2}\in\mathbb{R}^{d}$ are learnable parameters. For simplicity, we omitted 1D convolutions and for Mamba we wrote the matrices for the recursion of each row of ${\bm{H}}_{t}$ and set ${\bm{k}}_{t}=(k_{t,1},\dots,k_{t,n})^{\top}$ and ${\bm{W}}_{1}=({\bm{w}}_{1,1},\dots,{\bm{w}}_{1,n})^{\top}$.
${\bm{A}}({\bm{x}}_{t})$ ${\bm{B}}({\bm{x}}_{t})$ $\mathrm{dec}({\bm{H}}_{t},{\bm{x}}_{t})$ Mamba $\mathrm{Diag}\left(\exp\left(-\bm{\Delta}_{t}\exp({\bm{w}}_{1,i})\right)\right)$ $k_{t,i}\bm{\Delta}_{t}\odot{\bm{x}}_{t}$ $\psi({\bm{q}}_{t}^{\top}{\bm{H}}_{t}^{\top}+{\bm{w}}_{2}\odot{\bm{x}}_{t})$ GLA $\mathrm{Diag}\left(\bm{\alpha}_{t}\right)$ ${\bm{k}}_{t}{\bm{v}}_{t}^{\top}$ $\psi({\bm{q}}_{t}^{\top}{\bm{H}}_{t}^{\top})$ DeltaNet ${\bm{I}}-\beta_{t}{\bm{k}}_{t}{\bm{k}}_{t}^{\top}$ $\beta_{t}{\bm{k}}_{t}{\bm{v}}_{t}^{\top}$ $\psi({\bm{q}}_{t}^{\top}{\bm{H}}_{t}^{\top})$
@@ 4.1 Limitations of Current LRNNs
The parity $y_{t}\in\{0,1\}$ of a sequence of ones and zeros $x_{1}\dots x_{t}\in\{0,1\}^{t}$ is 1 if the total number of ones in the sequence is odd, and 0 if it’s even. Equivalent to addition modulo 2, it can be computed by summing the values in the input sequence and then applying the modulo 2 function: $y_{t}=(\sum_{i=1}^{t}x_{i})\bmod 2$. We can also express this as the linear recursion
@@
$h_{t}=a(x_{t})h_{t-1}+b(x_{t}),\quad h_{0}=0,\quad b(1)=a(0)=1,\quad a(1)=-1,\quad y_{t}=h_{t}.$ (3)
@@
A finite precision LRNN with finitely many layers as in (1) can solve parity for arbitrary input lengths, in particular, it can recognize the language $(11)^{*}$, only if in at least one layer, there exist ${\bm{x}}$ such that ${\bm{A}}({\bm{x}})$ has at least one eigenvalue $\lambda$ with $|\lambda|\geq 1$ and $\lambda\notin\{x\in\mathbb{R}\,:\,x\geq 0\}$.
The proof in Section B.1 uses the same core idea as the one in (Sarrof et al., 2024, Theorem 2). For one layer, we show that when ${\bm{x}}=1^{k}$ and the conditions for the eigenvalues of ${\bm{A}}(1)$ are not met, the mapping $k\mapsto{\bm{H}}_{k}$ and consequently also the one $k\mapsto\hat{\bm{y}}_{k}$ will be constant for large enough $k$ and in finite precision, while $k\mapsto y_{k}$, with $y_{k}$ being the parity of ${\bm{x}}$, alternates between 00 and $1$. To show this, we use the expression for the powers of the Jordan Canonical form of ${\bm{A}}(1)$, to prove that each element of ${\bm{A}}(1)^{k}$ either converges or diverges to a point in the complex infinity when $k\to\infty$.
We now study the problem of counting modulo $m$, which can be seen as an easier version of addition modulo $m$. For this problem, the input of length $k$ never changes and is equal to ${\bm{x}}=1^{k}$, while the correct output is $y_{k}=(\sum_{i=1}^{k}x_{i})\bmod m$.
The following theorem shows that to solve this problem, products of state-transition matrices must have at least one eigenvalue with a nonzero imaginary part and a modulus greater or equal to one.
@@
A finite precision LRNN with $L$ layers, each as in (1), can count modulo $m$, i.e. it can recognize the language $(1^{m})^{*}$, with $m$ not a power of two, only if there exist $i\in\{1,\dots,L\}$ and ${\bm{x}}_{1},\dots,{\bm{x}}_{2^{i-1}}$ such that for the $i$-th layer the product ${\bm{A}}({\bm{x}}_{1}){\bm{A}}({\bm{x}}_{2})\cdots{\bm{A}}({\bm{x}}_{2^{i-1}})$ has at least one eigenvalue $\lambda$ with $|\lambda|\geq 1$ and nonzero imaginary part.
The proof is in Section B.2. When $L=1$ a key step is to show that if ${\bm{A}}(1)$ has real (even negative) eigenvalues, the sequences $\{{\bm{H}}_{2k}\}_{k\in\mathbb{N}}$ and $\{{\bm{H}}_{2k+1}\}_{k\in\mathbb{N}}$ have a well defined limit. The proof for $L$ layers is done by induction using our assumption on the product of state-transition matrices.
@@ 4.2 Allowing Negative Eigenvalues
We focus on two classes of LRNNs determined by the structure of their state-transition matrices: diagonal (such as Mamba, Mamba2, and GLA) and generalized Householder (GH, as in DeltaNet) (i.e. non-diagonal). In particular, if we let ${\bm{s}}:\mathbb{R}^{l}\to[0,1]^{n}$, $\phi:\mathbb{R}^{l}\to[0,1]$ and ${\bm{v}}:\mathbb{R}^{l}\to\mathbb{R}^{n}$, being learnable functions such that $\left\lVert{\bm{v}}({\bm{x}})\right\rVert=1$ for every ${\bm{x}}\in\mathbb{R}^{l}$, then the state transition matrices of each layer of many LRNNs, such as those in Table 1, can be written as either
@@
where ${\bm{A}}_{\mathrm{diag}}({\bm{x}})$ is diagonal with eigenvalues $s({\bm{x}})_{i}\in[0,1]$, while ${\bm{A}}_{\mathrm{GH}}({\bm{x}})$ is GH with all eigenvalues equal to one except for the one associated to the eigenvector ${\bm{v}}({\bm{x}})$, which is equal to $1-\phi({\bm{x}})\in[0,1]$. To address the limitations discussed in the previous section, we propose the following modification that can be easily applied to any LRNN belonging to either class.
@@
We know from the previous section, that LRNNs with the modified state transition matrices can implement the solution to the parity problem by setting $s(1)=0$ and $\phi(1)=1$ so that if we consider a scalar recursion, then ${\bm{A}}^{-}_{\mathrm{diag}}(1)={\bm{A}}^{-}_{\mathrm{GH}}(1)=-1$. We have also shown that we cannot count modulo $3$ with diagonal state transition matrices, even when allowing negative eigenvalues. However, it is well known that counting modulo $m$ can be achieved by rotating a vector in $\mathbb{R}^{2}$ by an angle of $2\pi/m$ radians, and we can express a rotation matrix as a product of two reflection matrices, which are GH matrices with eigenvalues in $\{-1,1\}$. In other words, for any $m\in\mathbb{N}$ there exist unit norm vectors ${\bm{v}}_{1},{\bm{v}}_{2}\in\mathbb{R}^{2}$ such that
${\bm{R}}(\theta):=\left[\begin{array}[]{cc}\cos\theta&-\sin\theta\\ \sin\theta&\cos\theta\end{array}\right]=\left({\bm{I}}-2{\bm{v}}_{1}{\bm{v}}_{1}^{\top}\right)\left({\bm{I}}-2{\bm{v}}_{2}{\bm{v}}_{2}^{\top}\right),\quad\theta=\frac{2\pi}{m}.$
@@
Therefore, in the upcoming section, we examine the impact of our change to the eigenvalue range on state-transition matrices constructed as repeated products of GH matrices.

### v2

@@ 1 Introduction
Despite these successes, both Transformers and current LRNNs face a fundamental limitation: they struggle to learn how to track the state of even simple finite-state machines from sequences of state-transitions (Deletang et al., 2023). This limitation may impair performance on tasks such as entity tracking in narratives, handling nested structures in code, and other reasoning tasks that can benefit from maintaining and updating an internal state over time (Merrill et al., 2024). Even the simplest state-tracking task, computing the parity of a sequence of bits, cannot be solved by current LRNNs, while non-linear RNNs like LSTM (Hochreiter & Schmidhuber, 1997) and sLSTM (Beck et al., 2024) can track the state of any finite state machine. However, parallelizing non-linear RNNs across the sequence length presents significant challenges (Lim et al., 2024; Gonzalez et al., 2024).
Recently, Sarrof et al. (2024) demonstrated that the inability of diagonal LRNNs to solve the parity problem stems from the fact that the eigenvalues of their state-transition matrices are constrained to be positive. Specifically, they proved that finite precision diagonal LRNNs with exclusively positive real eigenvalues, cannot solve the parity problem for sequences of arbitrary length. However, their work did not provide empirical evidence showing that diagonal LRNNs with negative eigenvalues can be successfully trained to overcome this limitation. We prove that the same limitation also affects LRNNs with non-diagonal state-transition matrices, and further prove that complex eigenvalues are necessary to solve the more challenging task of modular counting (when the modulus is not a power of two). Our findings also apply to the GH matrices employed by DeltaNet, as they share the same eigenvalue limitations. To overcome this, we propose a simple yet powerful solution: extend the range of possible eigenvalues from $[0,1]$ to $[-1,1]$. This change enables state-tracking and significantly improves the expressivity of LRNNs without compromising their efficiency and training stability. As illustrated in Figure 1, it allows diagonal LRNNs to learn parity successfully. The code for part of our experiments is available at https://github.com/automl/unlocking_state_tracking
@@
    By extending the eigenvalue range, we significantly improve the state-tracking capabilities of LRNNs. We prove that LRNNs with state-transition matrices formed by products of generalized Householder (GH) matrices, each with eigenvalues in the range $[-1,1]$, can learn any regular language (Theorem 4), in some cases with just one layer (Theorem 3). Notably, this range extension allows LRNNs, using just one GH matrix (like DeltaNet), to learn substantially harder tasks, as the repeated composition of permutations of two elements, compared to diagonal LRNNs.
@@ 2 Related Work
Linear RNNs.  Linear RNNs encompass state-space models and causal, linear attention mechanisms. State-space models, originally used for continuous dynamical systems, inspired LRNN variants like S4 (Gu et al., 2022) and H4 (Fu et al., 2021) (see Tiezzi et al. (2024) for a survey). Recent advancements, such as Mamba (Gu & Dao, 2023; Dao & Gu, 2024), introduced input-dependent gating of the hidden state, significantly improving language modeling performance. Currently, linear attention has emerged as an alternative to classical softmax attention, with Katharopoulos et al. (2020) demonstrating that causal linear attention Transformers can be reformulated as RNNs with linear scaling in sequence length. Building on this, Yang et al. (2024a) proposed Gated Linear Attention (GLA), adding a gating mechanism similar to Mamba, while DeltaNet (Schlag et al., 2021; Yang et al., 2024b) and TTT-Linear (Sun et al., 2024) explored more expressive gating with non-diagonal state-transition matrices. Beck et al. (2024) recently proposed xLSTM, a successor LSTM (Hochreiter & Schmidhuber, 1997) which combines non-linear and linear RNNs.
@@
Expressivity Results.  Several studies have explored the expressive power of Transformers and RNNs (see e.g. (Merrill et al., 2020; Strobl et al., 2024; Bhattamishra et al., 2024)). Here, we focus on the ones most relevant to our work. While Hahn (2020) proved that Transformers cannot model periodic languages such as parity and some context-free languages at arbitrary sequence lengths, Liu et al. (2023) demonstrated that Transformers can learn shortcut solutions for solvable finite state automata, though these solutions lack generalizability to arbitrary sequence lengths and perform poorly out-of-distribution. Unlike RNNs, the high parallelizability of Transformers prevents them from learning unsolvable finite state automata (Merrill & Sabharwal, 2023). These findings typically use techniques from algebraic formal language theory (we refer to Liu et al. (2023) for a short tutorial) and circuit complexity, using the log-precision assumption and a number of layers scaling linearly or logarithmically with sequence length. While earlier research established Transformers’ Turing completeness, it relied on either arbitrary precision (Pérez et al., 2021) or arbitrary depth and weight sharing (Giannou et al., 2023). Diagonal LRNNs can simulate any RNN with infinite depth (Gu & Dao, 2023) and approximate regular enough functions when the state dimension grows linearly with sequence length (Orvieto et al., 2024). However, things change when depth and state size are fixed. Merrill et al. (2024) proved that finite-depth diagonal LRNNs, like Transformers, cannot learn unsolvable finite state automata when restricted to log-precision arithmetic. The work by Fan et al. (2024) highlights a similar limitation, while in a finite precision setting, Sarrof et al. (2024) showed that diagonal LRNNs with positive values in the state-transition matrix, while capable of learning all star-free languages, cannot solve even the simple parity problem, a non-star-free language recognizable by an automaton with two states. However, their analysis was limited to the diagonal case and they did not test the benefit of negative eigenvalues in practice. Unlike these works, we also study non-diagonal LRNNs that can still be trained efficiently at large scale. Irie et al. (2021; 2023) empirically show how state-tracking can be enabled by modifying DeltaNet as a fast weight programmer (Schmidhuber, 1992), but this makes its recurrence non-linear.
@@ 3.1 Linear Recurrent Neural Networks (LRNNs)
Table 1: Instances of LRNNs layers in (1), where $\bm{\alpha}_{t}{=}\mathrm{sigmoid}({\bm{W}}_{\alpha}{\bm{x}}_{t})$, $\bm{\Delta}_{t}{=}\mathrm{softplus}({\bm{W}}_{\Delta}{\bm{x}}_{t})$, $\beta_{t}{=}\mathrm{sigmoid}({\bm{w}}_{\beta}^{\top}{\bm{x}}_{t})$, while ${\bm{q}}_{t},{\bm{k}}_{t}\in\mathbb{R}^{n},{\bm{v}}_{t}\in\mathbb{R}^{d}$ are output of learnable functions of ${\bm{x}}_{t}$. Also, $\psi:\mathbb{R}^{d}\to\mathbb{R}^{d}$ is another learnable function usually containing an MLP and a normalization, while ${\bm{W}}_{1}\in\mathbb{R}^{n\times d}$, ${\bm{W}}_{\Delta}\in\mathbb{R}^{d\times l}$, ${\bm{W}}_{\alpha}\in\mathbb{R}^{n\times l}$, ${\bm{w}}_{\beta}\in\mathbb{R}^{l}$ and ${\bm{w}}_{2}\in\mathbb{R}^{d}$ are learnable parameters. For simplicity, we omitted 1D convolutions. For Mamba, the matrices in the first two columns represent the recurrence for each row of ${\bm{H}}_{t}$ and we set ${\bm{k}}_{t}{=}(k_{t,1},\dots,k_{t,n})^{\top}$, ${\bm{W}}_{1}{=}({\bm{w}}_{1,1},\dots,{\bm{w}}_{1,n})^{\top}$, $l=d$.
${\bm{A}}({\bm{x}}_{t})$ ${\bm{B}}({\bm{x}}_{t})$ $\mathrm{dec}({\bm{H}}_{t},{\bm{x}}_{t})$ Mamba $\mathrm{Diag}\left(\exp\left(-\bm{\Delta}_{t}\odot\exp({\bm{w}}_{1,i})\right)\right)$ $k_{t,i}\bm{\Delta}_{t}\odot{\bm{x}}_{t}$ $\psi({\bm{H}}_{t}^{\top}{\bm{q}}_{t}+{\bm{w}}_{2}\odot{\bm{x}}_{t})$ GLA $\mathrm{Diag}\left(\bm{\alpha}_{t}\right)$ ${\bm{k}}_{t}{\bm{v}}_{t}^{\top}$ $\psi({\bm{H}}_{t}^{\top}{\bm{q}}_{t})$ DeltaNet ${\bm{I}}-\beta_{t}{\bm{k}}_{t}{\bm{k}}_{t}^{\top}$ $\beta_{t}{\bm{k}}_{t}{\bm{v}}_{t}^{\top}$ $\psi({\bm{H}}_{t}^{\top}{\bm{q}}_{t})$
@@ 4.1 Limitations of Current LRNNs
The parity $y_{t}\in\{0,1\}$ of a sequence of ones and zeros $x_{1}\dots x_{t}\in\{0,1\}^{t}$ is 1 if the total number of ones in the sequence is odd, and 0 if it’s even. Equivalent to addition modulo 2, it can be computed by summing the values in the input sequence and then applying the modulo 2 function: $y_{t}=(\sum_{i=1}^{t}x_{i})\bmod 2$. We can also express this as the linear recurrence
@@
$h_{t}=a(x_{t})h_{t-1}+b(x_{t}),\quad h_{0}=b(0)=0,\quad b(1)=a(0)=1,\quad a(1)=-1,\quad y_{t}=h_{t}.$ (3)
@@
A finite precision LRNN with finitely many layers as in (1) can solve parity for arbitrary input lengths, in particular, it can recognize the language $(11)^{*}$, only if in at least one layer, there exist ${\bm{x}}$ such that ${\bm{A}}({\bm{x}})$ has at least one eigenvalue $\lambda\notin\{x\in\mathbb{R}\,:\,x\geq 0\}$.
The proof in Section B.1 uses the same core idea as the one in (Sarrof et al., 2024, Theorem 2). For one layer, we show that when ${\bm{x}}=1^{k}$ and the conditions for the eigenvalues of ${\bm{A}}(1)$ are not met, the mapping $k\mapsto{\bm{H}}_{k}$ and consequently also the one $k\mapsto\hat{\bm{y}}_{k}$ will be constant for large enough $k$ and in finite precision, while $k\mapsto y_{k}$, with $y_{k}$ being the parity of ${\bm{x}}$, alternates between 00 and $1$. To show this, we use the expression for the powers of the Jordan canonical form of ${\bm{A}}(1)$.
We now study the problem of counting modulo $m$, an easier version of addition modulo $m$ where the input of length $k$ never changes and is ${\bm{x}}=1^{k}$, while the correct output is $y_{k}=(\sum_{i=1}^{k}x_{i})\bmod m$. The following theorem shows that to solve this problem, products of state-transition matrices must have at least one eigenvalue with nonzero imaginary part.
@@
A finite precision LRNN with $L$ layers, each as in (1), can count modulo $m$, i.e. it can recognize the language $(1^{m})^{*}$, with $m$ not a power of two, only if there exist $i\in\{1,\dots,L\}$ and ${\bm{x}}_{1},\dots,{\bm{x}}_{2^{i-1}}$ such that for the $i$-th layer the product ${\bm{A}}({\bm{x}}_{1}){\bm{A}}({\bm{x}}_{2})\cdots{\bm{A}}({\bm{x}}_{2^{i-1}})$ has at least one eigenvalue $\lambda$ with nonzero imaginary part, i.e. $\lambda\notin\mathbb{R}$.
The proof is in Section B.2. When $L=1$ a key step is to show that if ${\bm{A}}(1)$ has real (even negative) eigenvalues, the sequences $\{{\bm{H}}_{2k}\}_{k\in\mathbb{N}}$ and $\{{\bm{H}}_{2k+1}\}_{k\in\mathbb{N}}$ have a well defined limit. The proof for $L>1$ layers is done by induction using our assumption on the product of state-transition matrices.
@@ 4.2 Allowing Negative Eigenvalues
We focus on two classes of LRNNs determined by the structure of their state-transition matrices: diagonal (such as Mamba, Mamba2, and GLA) and generalized Householder (GH, as in DeltaNet). In particular, if we let ${\bm{s}}:\mathbb{R}^{l}\to[0,1]^{n}$, $\phi:\mathbb{R}^{l}\to[0,1]$ and ${\bm{v}}:\mathbb{R}^{l}\to\mathbb{R}^{n}$, being learnable functions such that $\left\lVert{\bm{v}}({\bm{x}})\right\rVert=1$ for every ${\bm{x}}\in\mathbb{R}^{l}$, then the state transition matrices of each layer of many LRNNs, such as those in Table 1, can be written as either
@@
where ${\bm{A}}_{\mathrm{diag}}({\bm{x}})$ is diagonal with eigenvalues $s({\bm{x}})_{i}\in[0,1]$, while ${\bm{A}}_{\mathrm{GH}}({\bm{x}})$ is GH with all eigenvalues equal to one except for the one associated to the eigenvector ${\bm{v}}({\bm{x}})$, which is equal to $1-\phi({\bm{x}})\in[0,1]$. To address the limitations discussed in the previous section, we propose the following modification that can be easily applied to LRNNs belonging to either class.
@@
LRNNs with the modified state transition matrices can implement the solution to parity in (3) by setting $s(1)=0$ and $\phi(1)=1$ so that if we consider a scalar recursion, then ${\bm{A}}^{-}_{\mathrm{diag}}(1)=-1$. However, Theorem 2 shows that we cannot count modulo $3$ with diagonal state transition matrices, even when allowing negative eigenvalues. Despite this, it is well known that counting modulo $m$ can be achieved by rotating a vector in $\mathbb{R}^{2}$ by an angle of $2\pi/m$ radians, and we can express a rotation matrix as a product of two reflection matrices, which are GH matrices with eigenvalues in $\{-1,1\}$. In other words, for any $m\in\mathbb{N}$ there exist unit norm vectors ${\bm{v}}_{1},{\bm{v}}_{2}\in\mathbb{R}^{2}$ such that
${\bm{R}}(\theta):=\Big{[}\begin{array}[]{cc}\cos\theta&-\sin\theta\\ \sin\theta&\cos\theta\end{array}\Big{]}=\left({\bm{I}}-2{\bm{v}}_{1}{\bm{v}}_{1}^{\top}\right)\left({\bm{I}}-2{\bm{v}}_{2}{\bm{v}}_{2}^{\top}\right),\quad\theta=\frac{2\pi}{m}.$
@@
Therefore, in the next section, we examine the impact of our change to the eigenvalue range on state-transition matrices constructed as repeated products of GH matrices.